OpenStudy (anonymous):

Solve each system of equations using the method of your choice: Substitution, Elimination, or Graphing. y = x + 1 2x + y = 10 Select one: (1, 5) (3, 4) (2, 4) (4, 2) Solve each system of equations using the method of your choice: Substitution, Elimination, or Graphing. 3x + 5y = 2 x + 4y = -4 Select one: (3, 3) (4, -2) (2, 4) (0, 4) Tricia has $15 and she earns $6 an hour walking dogs. Mike has $9 and he earns $8 an hour walking dogs. Use the system of linear equations to find the number of hours of work after which Tricia and Mike will have the same amount of money.

OpenStudy (anonymous):

y = x + 1 2x + y = 10 with this problem, I'd use substitution. Substitute the 1st equation into the second. You'll have: 2x+(x+1)=10 Now solve for x

OpenStudy (anonymous):

hm?

OpenStudy (anonymous):

I am a little confused on this one

OpenStudy (anonymous):

I'm doing it on paper now.

OpenStudy (anonymous):

I substituted in x+1 where I saw a y in the second problem. 2x+(x+1)=10

OpenStudy (anonymous):

When I simplify (add like terms) I get 3x+1=10

OpenStudy (anonymous):

Well I did it and what I got wasn't in one of my options..

OpenStudy (anonymous):

3x+1=10 3x+1-1=10-1 3x=9 x=3

OpenStudy (anonymous):

Yes thats what I did

OpenStudy (anonymous):

(1, 5) (3, 4) (2, 4) (4, 2)

OpenStudy (anonymous):

I gotta choose out of those

OpenStudy (anonymous):

OK, to get y, substitute back into the original equations 3+1=y 4=y (3,4)

OpenStudy (anonymous):

second choice

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

3x + 5y = 2 x + 4y = -4 You can solve this by substitution if you solve the 2nd equation for x...then solve for y...or you could use elimination. Since I've done a substitution problem, let's use elimination. I'm going to eliminate the x's.

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

3x + 5y = 2 3x + 5y = 2 3x+5y=2 x + 4y = -4 -3(x +4y=-4) -3x-12y=12 --------- -7y=14 y=-2 Now solve for x

OpenStudy (anonymous):

OK one second

OpenStudy (anonymous):

Ok i got (4,-2)

OpenStudy (anonymous):

Right...great!

OpenStudy (anonymous):

Yay!

OpenStudy (anonymous):

Tricia has $15 and she earns $6 an hour walking dogs. Mike has $9 and he earns $8 an hour walking dogs. Use the system of linear equations to find the number of hours of work after which Tricia and Mike will have the same amount of money. m = 6h + 15 m = 8h + 9 Select one: 7 hours 5 hours 1 hour 3 hours

OpenStudy (anonymous):

Could you help me with that one?

OpenStudy (anonymous):

sure...since both equations are m=something, I'd put each equation equal to each other. 6h+15=8h+9 solve for h

OpenStudy (anonymous):

I'm doing it

OpenStudy (anonymous):

I sub. 9 from both sides as step one and now I have 6h + 6 = 8h?

OpenStudy (anonymous):

then I subtracted 6h from each side and I got 6 = 2h and divided two from each side and got 3=h

OpenStudy (anonymous):

yes...so after 3 hours, they're making the same amount of money

OpenStudy (anonymous):

good job

OpenStudy (anonymous):

Ok :) Thank you

OpenStudy (anonymous):

Have a great day!

OpenStudy (anonymous):

You too